# ELLA: Efficient Lifelong Learning for Adapters in Large Language Models

### v1

@@ 4.5 Results
On the Standard CL benchmark, ELLA achieves an average accuracy of $79.9$ on T55-Large, outperforming the strongest replay-free competitor qiao2024learn and exceeding DATA liao2025data and Recurrent KIF feng2025recurrent, both of which depend on data rehearsal. On the more demanding Long Sequence (LS) setup, ELLA improves over qiao2024learn; du2024unlocking by upto $4.3\%$, demonstrating robustness under long horizons where interference typically compounds. This is further examined in Sec. 5 and Fig. 5. On TRACE, which stresses cross-domain CL adaptation, ELLA achieves a dramatic gain of up to +23.3+23.3, showing its ability to generalize to reasoning, code, and multilingual tasks.
Beyond average accuracy, Fig. 2 and 3 reveal deeper dynamics. ELLA achieves the best BWT, sharply reducing forgetting, while also obtaining the highest FWT, signaling effective knowledge reuse across tasks. This is crucial: unlike orthogonality-based approaches wang2023orthogonal; qiao2024learn, which guarantee stability but block positive transfer, ELLA explicitly allows low-energy subspace reuse, yielding consistent forward gains. Conversely, replay-based approaches achieve modest transfer but at the cost of large buffers and compute. Notably, ELLA sustains high performance on interference-sensitive tasks like QQP, IMDB, and DBpedia, where baselines degrade sharply. ELLA thus demonstrates that selective subspace decorrelation provides a more principled way to reconcile stability and plasticity than existing strategies.
Equally important, ELLA achieves these gains with no replay buffer, no generative augmentation, and no task identity signals. Replay-heavy methods liao2025data; feng2025recurrent not only incur large memory overheads but also encounter out-of-memory failures on LLaMA-3.1 8B when trained on GPUs smaller than H100100s. ELLA avoids these pitfalls entirely, remaining lightweight, scalable, and stable even at billion-parameter scale.
@@ 4.6 Task Generalization
LLMs that are continually trained on new tasks often exhibit a decline in general performance, revealing catastrophic forgetting of their original capabilities. To assess this aspect, we evaluate the generalization ability of ELLA in the context of continual learning, focusing on its performance on five unseen cross-task benchmarks: MMLU hendrycks2020measuring, GSM88cobbe2021training, BBH suzgun2022challenging, AGIEval zhong2023agieval, and PIQA bisk2020piqa. More details on these datasets are provided in Appendix B.
@@ B.1 Datasets
Train Tasks. Tables 5 and  6 provide detailed information on the datasets utilized in our continual learning (CL) experiments. Table 5 presents the $15$ datasets included in the Long Sequence Benchmark razdaibiedina2023progressive, while Table 6 outlines the $8$ datasets from TRACE wang2023trace. Both tables include the corresponding evaluation metrics for each dataset.
Unseen tasks. We select the following datasets to test generalization performance post continual adaptation: (1) Multitask Language Understanding (MMLU) hendrycks2020measuring, which includes multiple-choice questions across $57$ subjects. (2) GSM8K cobbe2021training, which is a high-quality linguistically diverse multi-step elementary math reasoning dataset. (3) BIG-Bench Hard (BBH) suzgun2022challenging, which includes $27$ challenging tasks spanning arithmetic, symbolic reasoning, and more, derived from BIG-Bench (BB) srivastava2022beyond. Most of the data consists of multiple-choice questions. (4) AGIEval zhong2023agieval, which includes a wide range of high-quality official entrance exams, qualifying exams, and advanced competitions tailored to human participants. (5) PIQA bisk2020piqa which is a dataset for commonsense reasoning, and was created to investigate the physical knowledge of existing models in NLP.
@@
Table 5: The details of 15 classification datasets in the Long Sequence Benchmark razdaibiedina2023progressive. The first five tasks correspond to the standard CL benchmark zhang2015character.
@@
Table 6: The overview of dataset statistics in TRACE wang2023trace. The ‘Source’ indicates the origin of the context. ‘Avg len’ denotes the average length, measured in word count for English, German, and code datasets, and in character count for Chinese datasets. ‘SARI’ is a score that is specific to evaluating simplification tasks.
@@
Table 7: Seven distinct orders of task sequences were employed for the experiments in continual learning. Orders 1-3 align with the Standard CL Benchmarks, as adopted in previous studies liao2025data. Orders 4-6 pertain to the Long Sequence Benchmarks, which encompass a total of 15 tasks razdaibiedina2023progressive. Order 7 refers to the TRACE benchmark, specifically designed for LLMs, and comprises eight datasets wang2023trace.
@@ B.3 Baselines
We compare our method against a comprehensive set of recent CL baselines, detailed as follows: (1) SeqFT sequentially fine-tunes all model parameters without CL mechanisms like regularization or replay (RorR). (2) SeqLoRA applies fixed-size LoRA tuning per task with/without replay. (3) IncLoRA allows incremental learning of new LoRA parameters for each task without constraints. (4) SeqSVD uses fixed-rank SVD adapters trained sequentially without RorR. (5) EWC kirkpatrick2017overcoming finetunes LoRA with a regularization loss to prevent interference with past tasks. (6) LwF li2017learning distills the model of the last task using the current task data. (7) L2P wang2022learning instantiates a prompt pool for adaptive prompt selection and prompt tuning for individual samples. (8) LFPT5 qin2021lfpt5 learns soft prompts and generates pseudo-data for replay. (9) L-CL trains SVD adapters incrementally with SVD regularization. (10) B-CL applies SVD regularization with gradient projection. (11) O-LoRA wang2023orthogonal enforces orthogonality between LoRA updates across tasks. (12)  MIGU du2024unlocking updates parameters based on gradient magnitude. (13) LB-CL qiao2024learn initializes low-rank matrix parameters in new tasks from specific past parameters besides enforcing orthogonality. (14) DATA liao2025data decomposes attention into high and low rank subspaces, and leverages orthogonality with optional replay. (15) Recurrent KIF feng2025recurrent uses an expensive recurrent mechanism to modulate task-specific adapter reuse using replay.
@@
Instruction-following is a fundamental capability for LLMs to serve as an effective interface between humans and AI systems wei2021finetuned; ouyang2022training. We adopt instruction tuning as our training paradigm for two key reasons: (1) it enables the incorporation of human prior knowledge via explicit instructions, thereby facilitating more efficient learning; and (2) it enhances generalization by guiding the model to learn underlying principles that apply across tasks.
@@ B.6 Implementation Details
Our implementation for ELLA uses PyTorch v2.0.1 paszke2019pytorch and Deepspeed v0.10.0 rasley2020deepspeed. Our experiments were conducted on machines equipped with 44 A4040 GPUs/ 88 V100100 GPUs for ELLA and all baselines, except DATA/DATA+Replay, which required 22 H100100 GPUs. For all orders of task streams for the Standard-CL Benchmark and the Long Sequence Benchmark, we trained the models with one epoch using the AdamW optimizer loshchilov2017decoupled and WarmupLR scheduler deepspeed_schedulers with a total batch size of $32$. We used a constant learning rate of 1e31e-3 for T55 and 1e41e-4 for LLaMA, a dropout rate of $0.1$, and a weight decay rate of 0. For TRACE Order $7$ (C-STANCE, FOMC, MeetingBank, Py150, ScienceQA, NumGLUE-cm, NumGLUE-ds, 20Minuten), we trained with $5000$ samples $5$, 33, 77, 55, 33, 55, 55, 77 epochs respectively. The coefficient $\lambda$ for different task orders has been reported in Table 9. $\lambda$ was selected based on performance on a small, held-out validation set. To establish an effective search range, we followed the common heuristic of choosing values that scale the regularization term, $\mathcal{L}_{\text{ELLA}}$ (promoting stability), to a similar order of magnitude as the current task accuracy loss (promoting plasticity) for balanced learning. We observed that performance was robust across this range and that a single $\lambda$ often generalized well across multiple subsequent tasks. The LoRA rank was set to $8$ for all experiments, and the proportion of past task data mixed in for replay methods was set to $2\%$ of the original training set.

### v2

@@ 4.5 Results
On the Standard CL benchmark, ELLA achieves an average accuracy of $79.9$ on T55-Large, outperforming the strongest replay-free competitor Qiao and Mahdavi (2024) and exceeding DATA Liao et al. (2025) and Recurrent KIF Feng et al. (2025), both of which depend on data rehearsal. On the more demanding Long Sequence (LS) setup, ELLA improves over Qiao and Mahdavi (2024); Du et al. (2024) by upto $4.3\%$, demonstrating robustness under long horizons where interference typically compounds. This is further examined in Sec. 5 and Fig. 5. On TRACE, which stresses cross-domain CL adaptation, ELLA achieves a dramatic gain of up to +23.3+23.3, showing its ability to generalize to reasoning, code, and multilingual tasks.
Beyond average accuracy, Fig. 2 and 3 reveal deeper dynamics. ELLA achieves the best BWT, sharply reducing forgetting, while also obtaining the highest FWT, signaling effective knowledge reuse across tasks. This is crucial: unlike orthogonality-based approaches Wang et al. (2023a); Qiao and Mahdavi (2024), which guarantee stability but block positive transfer, ELLA explicitly allows low-energy subspace reuse, yielding consistent forward gains. Conversely, replay-based approaches achieve modest transfer but at the cost of large buffers and compute. Notably, ELLA sustains high performance on interference-sensitive tasks like QQP, IMDB, and DBpedia, where baselines degrade sharply. ELLA thus demonstrates that selective subspace decorrelation provides a more principled way to reconcile stability and plasticity than existing strategies.
Equally important, ELLA achieves these gains with no replay buffer, no generative augmentation, and no task identity signals. Replay-heavy methods Liao et al. (2025); Feng et al. (2025) not only incur large memory overheads but also encounter out-of-memory failures on LLaMA-3.1 8B when trained on GPUs smaller than H100100s. ELLA avoids these pitfalls entirely, remaining lightweight, scalable, and stable even at billion-parameter scale.
@@ 4.6 Task Generalization
LLMs that are continually trained on new tasks often exhibit a decline in general performance, revealing catastrophic forgetting of their original capabilities. To assess this aspect, we evaluate the generalization ability of ELLA in the context of continual learning, focusing on its performance on five unseen cross-task benchmarks: MMLU Hendrycks et al. (2020), GSM88Cobbe et al. (2021), BBH Suzgun et al. (2022), AGIEval Zhong et al. (2023), and PIQA Bisk et al. (2020). More details on these datasets are provided in Appendix B.
@@ B.1 Datasets
Train Tasks. Tables 5 and  6 provide detailed information on the datasets utilized in our continual learning (CL) experiments. Table 5 presents the $15$ datasets included in the Long Sequence Benchmark Razdaibiedina et al. (2023), while Table 6 outlines the $8$ datasets from TRACE Wang et al. (2023b). Both tables include the corresponding evaluation metrics for each dataset.
Unseen tasks. We select the following datasets to test generalization performance post continual adaptation: (1) Multitask Language Understanding (MMLU) Hendrycks et al. (2020), which includes multiple-choice questions across $57$ subjects. (2) GSM8K Cobbe et al. (2021), which is a high-quality linguistically diverse multi-step elementary math reasoning dataset. (3) BIG-Bench Hard (BBH) Suzgun et al. (2022), which includes $27$ challenging tasks spanning arithmetic, symbolic reasoning, and more, derived from BIG-Bench (BB) Srivastava et al. (2022). Most of the data consists of multiple-choice questions. (4) AGIEval Zhong et al. (2023), which includes a wide range of high-quality official entrance exams, qualifying exams, and advanced competitions tailored to human participants. (5) PIQA Bisk et al. (2020) which is a dataset for commonsense reasoning, and was created to investigate the physical knowledge of existing models in NLP.
@@
Table 5: The details of 15 classification datasets in the Long Sequence Benchmark Razdaibiedina et al. (2023). The first five tasks correspond to the standard CL benchmark Zhang et al. (2015).
@@
Table 6: The overview of dataset statistics in TRACE Wang et al. (2023b). The ‘Source’ indicates the origin of the context. ‘Avg len’ denotes the average length, measured in word count for English, German, and code datasets, and in character count for Chinese datasets. ‘SARI’ is a score that is specific to evaluating simplification tasks.
@@
Table 7: Seven distinct orders of task sequences were employed for the experiments in continual learning. Orders 1-3 align with the Standard CL Benchmarks, as adopted in previous studies Liao et al. (2025). Orders 4-6 pertain to the Long Sequence Benchmarks, which encompass a total of 15 tasks Razdaibiedina et al. (2023). Order 7 refers to the TRACE benchmark, specifically designed for LLMs, and comprises eight datasets Wang et al. (2023b).
@@ B.3 Baselines
We compare our method against a comprehensive set of recent CL baselines, detailed as follows: (1) SeqFT sequentially fine-tunes all model parameters without CL mechanisms like regularization or replay (RorR). (2) SeqLoRA applies fixed-size LoRA tuning per task with/without replay. (3) IncLoRA allows incremental learning of new LoRA parameters for each task without constraints. (4) SeqSVD uses fixed-rank SVD adapters trained sequentially without RorR. (5) EWC Kirkpatrick et al. (2017) finetunes LoRA with a regularization loss to prevent interference with past tasks. (6) LwF Li and Hoiem (2017) distills the model of the last task using the current task data. (7) L2P Wang et al. (2022b) instantiates a prompt pool for adaptive prompt selection and prompt tuning for individual samples. (8) LFPT5 Qin and Joty (2021) learns soft prompts and generates pseudo-data for replay. (9) L-CL trains SVD adapters incrementally with SVD regularization. (10) B-CL applies SVD regularization with gradient projection. (11) O-LoRA Wang et al. (2023a) enforces orthogonality between LoRA updates across tasks. (12)  MIGU Du et al. (2024) updates parameters based on gradient magnitude. (13) LB-CL Qiao and Mahdavi (2024) initializes low-rank matrix parameters in new tasks from specific past parameters besides enforcing orthogonality. (14) DATA Liao et al. (2025) decomposes attention into high and low rank subspaces, and leverages orthogonality with optional replay. (15) Recurrent KIF Feng et al. (2025) uses an expensive recurrent mechanism to modulate task-specific adapter reuse using replay.
@@
Instruction-following is a fundamental capability for LLMs to serve as an effective interface between humans and AI systems Wei et al. (2021); Ouyang et al. (2022). We adopt instruction tuning as our training paradigm for two key reasons: (1) it enables the incorporation of human prior knowledge via explicit instructions, thereby facilitating more efficient learning; and (2) it enhances generalization by guiding the model to learn underlying principles that apply across tasks.
@@ B.6 Implementation Details
Our implementation for ELLA uses PyTorch v2.0.1 Paszke et al. (2019) and Deepspeed v0.10.0 Rasley et al. (2020). Our experiments were conducted on machines equipped with 44 A4040 GPUs/ 88 V100100 GPUs for ELLA and all baselines, except DATA/DATA+Replay, which required 22 H100100 GPUs. For all orders of task streams for the Standard-CL Benchmark and the Long Sequence Benchmark, we trained the models with one epoch using the AdamW optimizer Loshchilov and Hutter (2017) and WarmupLR scheduler 10 with a total batch size of $32$. We used a constant learning rate of 1e31e-3 for T55 and 1e41e-4 for LLaMA, a dropout rate of $0.1$, and a weight decay rate of 0. For TRACE Order $7$ (C-STANCE, FOMC, MeetingBank, Py150, ScienceQA, NumGLUE-cm, NumGLUE-ds, 20Minuten), we trained with $5000$ samples $5$, 33, 77, 55, 33, 55, 55, 77 epochs respectively. The coefficient $\lambda$ for different task orders has been reported in Table 9. $\lambda$ was selected based on performance on a small, held-out validation set. To establish an effective search range, we followed the common heuristic of choosing values that scale the regularization term, $\mathcal{L}_{\text{ELLA}}$ (promoting stability), to a similar order of magnitude as the current task accuracy loss (promoting plasticity) for balanced learning. We observed that performance was robust across this range and that a single $\lambda$ often generalized well across multiple subsequent tasks. The LoRA rank was set to $8$ for all experiments, and the proportion of past task data mixed in for replay methods was set to $2\%$ of the original training set.